# Clear Nights Ahead: Towards Multi-Weather Nighttime Image Restoration

### v1

@@ Abstract
Restoring nighttime images affected by multiple adverse weather conditions is a practical yet under-explored research problem, as multiple weather conditions often coexist in the real world alongside various lighting effects at night. This paper first explores the challenging multi-weather nighttime image restoration task, where various types of weather degradations are intertwined with flare effects. To support the research, we contribute the AllWeatherNight dataset, featuring large-scale high-quality nighttime images with diverse compositional degradations, synthesized using our introduced illumination-aware degradation generation. Moreover, we present ClearNight, a unified nighttime image restoration framework, which effectively removes complex degradations in one go. Specifically, ClearNight extracts Retinex-based dual priors and explicitly guides the network to focus on uneven illumination regions and intrinsic texture contents respectively, thereby enhancing restoration effectiveness in nighttime scenarios. In order to better represent the common and unique characters of multiple weather degradations, we introduce a weather-aware dynamic specific-commonality collaboration method, which identifies weather degradations and adaptively selects optimal candidate units associated with specific weather types. Our ClearNight achieves state-of-the-art performance on both synthetic and real-world images. Comprehensive ablation experiments validate the necessity of AllWeatherNight dataset as well as the effectiveness of ClearNight. Project page: https://henlyta.github.io/ClearNight/mainpage.html.
1 Introduction
(a) ClearNight is the first unified framework that effectively removes multiple weather degradations within nighttime images in one go
(b) benchmark results across various types of weather degradations, where rain and snow scenes contain images affected by compound weather degradations, and ClearNight achieves the best average results over all images
Figure 1: Comparison of our introduced ClearNight with state-of-the-art adverse weather image restoration approaches. ClearNight exhibits robust generalization across diverse weather conditions and shows superior effectiveness in handling mixed degradations.
@@ 1 Introduction
Table 1: Comparison to related adverse weather datasets. H, RS, RD and S indicate haze, rain streak, raindrop and snow, respectively. P refers to the presence of flares and normal lighting image pairs in datasets. Syn. denotes synthesized images.
Datasets H RS RD S P # data Syn./Real Day/Night
Outdoor-Rain [30] 10.5K Syn. Day
RainDS [40] 5.8K Syn.&Real Day
BID-II [13] 32.6K Syn. Day
WeatherStream [62] 202K Real Day
CDD-11 [12] 15K Syn. Day
Raindrop Clarity[21] 15.2K Real Day&Night
Night&YellowHaze [31] 43.8K Syn. Night
NHC&NHM&NHR [64] 17K Syn.&Real Night
GTA5[52] 0.8K Syn. Night
UNREAL-NH [33] 10.3K Syn.&Real Night
GTAV-NightRain [59] 12.8K Syn. Night
GTAV-NightRain-H [60] 5.5K Syn. Night
RVSD [3] 101.9K Syn. Night
AllWeatherNight (ours) 10K Syn.&Real Night
Figure 2: Real-world and synthetic samples in AllWeatherNight dataset. The synthetic images effectively simulate real-world nighttime scenes with various realistic adverse weather and flare degradations.
@@
Image restoration under adverse weather conditions is a vital preprocessing step in many computer vision applications, such as autonomous driving [17, 50, 51, 56] and video surveillance [44, 14, 35]. Although significant efforts have been made to tackle the issues posed by adverse weather images, prior works [6, 55, 45, 53, 58] often neglect the complexities of lighting conditions and flares, limiting their effectiveness in nighttime scenarios.
Unlike daytime scenes, adverse weather degradations and lighting are closely intertwined in real nighttime scenes, severely obscuring the background contents. Although several works [64, 33, 60] have explored nighttime adverse weather image restoration, they typically focus only on a single type of weather condition, leading to poor performance in real-world images with complex scenarios (see Figure 1). Since multiple severe weather conditions usually co-occur in the real-world nighttime scenes, restoring images affected by multiple degradations is practical yet remains under-explored. This motivates us to investigate a unified framework for nighttime image restoration that can effectively handle multiple weather degradations in one go, while also mitigating the complex interactions between uneven lighting and adverse weather effects.
@@
However, to the best of our knowledge, there is no existing dataset that is applicable to multi-weather nighttime image restoration. As summarized in Table 1, all previous datasets [12, 21, 33, 60, 3], overlook nighttime scenarios with multiple adverse weather conditions under non-uniform illumination. To facilitate the research, we construct AllWeatherNight, a dataset for restoring nighttime images affected by multiple adverse weather effects and flares. Constructing such a dataset is challenging due to the scarcity of high-quality degraded-clean image pairs.
We design an illumination-aware degradation generation method to emulate real-world scenarios, which synthesizes the combined effect of uneven lighting and diverse adverse weather. The images generated with our illumination-aware synthesis not only achieve higher human perceptual preferences but also lead to improved performance on real-world images when used for training. In total, we generate 9K highly realistic synthetic image pairs for model training and evaluation (see Figure 2). In addition, we meticulously collect 1K nighttime photos to further assess models’ effectiveness and generalization capabilities on real images.
In nighttime scenes, the interplay between artificial lighting and multiple adverse weather effects complicates image restoration, as their intertwined visual manifestations make it challenging for existing unified models [6, 68, 65, 48] to effectively remove diverse degradations and lead to suboptimal performance. To address these issues, we present ClearNight, the first unified framework tailored for multi-weather nighttime image restoration, integrating the Retinex-based dual prior guidance as well as the weather-aware dynamic specificity and commonality collaboration.
@@
Specifically, we extract multi-scale illumination and reflectance priors based on Retinex theory [10, 27], for explicit disentanglement of lighting and texture information within degraded images. The illumination prior guides the model to focus on illumination-related regions, enabling more effective restoration under uneven lighting conditions. The reflectance prior enhances the feature representation of various texture information, mitigating weather-induced degradations and recovering clear background details.
@@
In addition, a specificity-commonality dual-branched architecture is devised to better represent the complex degradations caused by multiple types of adverse weather conditions, where the specificity branch is dynamically constructed and synergizes with the commonality branch based on the different weather types present in the input images. The dynamic specificity branch contains diverse sub-networks, each formed by selecting different combinations of candidate units [41, 66, 19]. To further enhance the capabilities of handling multi-weather degradations, we try to associate these units with specific weather types. Concretely, an auxiliary weather instructor is introduced to identify weather degradations and implicitly guide the dynamic allocator. Thereby, the dynamic allocator becomes weather-aware and is encouraged to select consistent units for the same types of adverse weather conditions.
@@
Figure 3: Visualization of four synthesized variants of adverse weather nighttime images, where Weather Only and Flare Only denote synthesis with illumination-aware weather degradation and flare, respectively. Ours involves synthesis with both types of degradations.
@@
  • We propose a unified framework to recover multi-weather nighttime images, which integrates the Retinex-based dual prior guidance as well as the weather-aware dynamic specificity-commonality collaboration method, tailored to address challenging uneven lighting conditions and diverse weather degradations entangled in nighttime scenes.
@@
  • Our approach effectively removes various types of degradations and cleans nighttime images. It outperforms state-of-the-art methods on both synthetic and real-world adverse weather nighttime images by a non-trivial margin.
2 Related Work
@@
Adverse Weather Nighttime Image Restoration. Early works [39, 63] typically relied on physical priors and statistical assumptions, limiting their effectiveness and robustness in handling complex real-world scenarios. To address the limitations, several data-driven methods have been developed, achieving impressive results in restoring nighttime images degraded by haze [31, 52, 64, 20, 33, 9], rain [60] and snow [3]. However, existing works focus on single-type degradations and overlook the fact that real-world nighttime scenes often involve multiple simultaneous degradations caused by adverse weather and uneven lighting.
@@
Multi-Weather Image Restoration. Multi-weather image restoration aims to restore various weather-degraded scenes using a single model [6, 26, 68, 24, 53, 8, 1, 12]. Recently, numerous Transformer-based approaches are developed, which investigate weather type queries [23], adverse weather pixel categorization [46], texture-guided appearance flow [47], codebook prior fusion [55], and etc. Moreover, several recent works [69, 4, 65] leverage remarkable generative capabilities of diffusion models for restoration performance boosting. To the best of our knowledge, all the previous methods focus on daytime scenes and ignore the entanglement of multiple adverse weather conditions with non-uniform illumination in nighttime scenes.
@@
Adverse Weather Datasets. Most previous datasets [49, 11, 54, 30, 40, 13, 62] only focuses on adverse weather restoration in daytime scenarios. A few recent studies explore this task in nighttime scenarios and build datasets for haze [31, 64, 33], rain [59, 60] and snow scenes [3]. However, each existing nighttime dataset only considers a single type of weather condition, ignoring the challenge of mixed weather conditions commonly found in the real world.
3 AllWeatherNight Dataset
For advancing the multi-weather nighttime image restoration, we construct AllWeatherNight, a new dataset featuring nighttime images with multiple adverse weather and uneven lighting conditions. We believe that handling mixed weather conditions is more practical for real-world scenarios, since multiple weather phenomena often co-occur, each causing unique distortions. Furthermore, flares in real-world photos obscure important details and distort the background, especially when intertwined with weather degradations (see Figure 2). This interplay significantly complicates image restoration, challenging models to maintain fidelity and enhance visibility in nighttime scenes.
3.1 Dataset Curation
Data Collection and Filtering. We repurpose images from a large-scale driving video dataset [56], which offers geographic and environmental diversity, and from a low-light object detection dataset [36] covering scenes from very dark to twilight, using them as uncorrupted ground truths. We empirically design several rules to first filter out irrelevant daytime images and unsuitable nighttime images. We use average brightness, average gradient and grayscale variance product metrics to identify and select nighttime images with adequate visibility, rich textures and high clarity. After initial data filtering, we manually select 1,000 diverse images from each dataset as ground-truths. We also crawl and select 1,000 real-world nighttime images with various adverse weather conditions from the Internet and existing datasets [29, 49, 34, 61, 11, 20] to evaluate the model performance on real-world data.
(a) human $\downarrow$ on syn. images
@@ 3.1 Dataset Curation
Figure 4: Comparison of 4 different synthesized variants of nighttime images, where W and F indicate synthesis with illumination-aware weather degradation and flare effect respectively.
Illumination-Aware Degradation Generation. We observe that uneven lighting conditions in real-world nighttime scenes often interact with adverse weather degradations, yet this phenomenon is largely overlooked by prior works [30, 64, 60]. To synthesize more realistic nighttime images with adverse weather conditions, we introduce an illumination-aware degradation generation approach.
Given a clean nighttime image $\mathbf{X}$, the illumination-aware weather-degraded image $\mathbf{X}_{\textit{d}}$ is synthesized by
$\mathbf{X}_{\textit{d}}=\mathbf{X}_{\textit{flare}}+\sum_{e\in\mathcal{E}}\mathbbm{1}_{e}\cdot\left(\omega_{e}\cdot\mathrm{G}_{e}(\mathbf{X}_{\textit{flare}})\right),\vspace{-0.5em}$ (1)

### v2

@@ Abstract
Restoring nighttime images affected by multiple adverse weather conditions is a practical yet under-explored research problem, as multiple weather degradations usually coexist in the real world alongside various lighting effects at night. This paper first explores the challenging multi-weather nighttime image restoration task, where various types of weather degradations are intertwined with flare effects. To support the research, we contribute the AllWeatherNight dataset, featuring large-scale nighttime images with diverse compositional degradations. By employing illumination-aware degradation generation, our dataset significantly enhances the realism of synthetic degradations in nighttime scenes, providing a more reliable benchmark for model training and evaluation. Additionally, we propose ClearNight, a unified nighttime image restoration framework, which effectively removes complex degradations in one go. Specifically, ClearNight extracts Retinex-based dual priors and explicitly guides the network to focus on uneven illumination regions and intrinsic texture contents respectively, thereby enhancing restoration effectiveness in nighttime scenarios. Moreover, to more effectively model the common and unique characteristics of multiple weather degradations, ClearNight performs weather-aware dynamic specificity and commonality collaboration that adaptively allocates optimal sub-networks associated with specific weather types. Comprehensive experiments on both synthetic and real-world images demonstrate the necessity of the AllWeatherNight dataset and the superior performance of ClearNight.
Project Page:https://henlyta.github.io/ClearNight/
Introduction
Figure 1: ClearNight is the first multi-weather nighttime image restoration framework, which effectively removes complex and coupled weather and flare degradations in one go.
[…]
Outdoor-Rain
RainDS
BID-II
WeatherStream
CDD-11
Raindrop Clarity
Night/YellowHaze
NHC/NHM/NHR
GTA5
UNREAL-NH
GTAV-NightRain
RVSD
AllWeatherNight
@@ Introduction
Table 1: Comparison to related adverse weather datasets. H, RS, RD and S indicate haze, rain streak, raindrop and snow, respectively. F denotes the presence of synthesized flare images. Syn denotes synthesized images. Night denotes the presence of nighttime data.
@@
Image restoration under adverse weather conditions is a vital preprocessing step in many computer vision applications, such as autonomous driving (BDD100K; quan2021holistic; driving2; xu2024mc; xu2022h2fa) and video surveillance (liu2023generating; li2024; li2025; AAAI2026; wu2025bvinet). Although significant efforts have been made to tackle the issues posed by adverse weather images, prior works (HECCV2024; LCVPR2024; GOUBICML2024) often neglect the complexities of lighting conditions (e.g., flare effects) and their intricate interplay with weather degradations, limiting the effectiveness in real-world applications.
In particular, nighttime scenes exacerbate such problem, where adverse weather degradations and uneven illumination are tightly coupled, severely obscuring background contents. Although a few works (NACM2020; NACM20232; NICCV2023) have explored nighttime adverse weather image restoration, they seldom take into account these unique characteristics. More critically, existing methods are limited to handling a single type of degradations, which leads to unsatisfactory performance in complex real-world scenarios where multiple adverse weather conditions frequently co-occur (see Fig. 1). To better meet the demands of real-world applications, this paper first explores a highly practical yet remains largely under-explored task: multi-weather nighttime image restoration. In this context, two critical challenges must be addressed: ❶ the scarcity of realistic multi-weather training samples and ❷ the insufficient ability of existing models to effectively address entangled degradations in nighttime scenes.
@@
To the best of our knowledge, there is no existing dataset that is applicable to multi-weather nighttime image restoration. As summarized in Tab. 1, all previous datasets (OneRestoreECCV2024; NRaindropECCV2024; NACM20232; NICCV2023; SICCV2023) overlook nighttime scenarios with multiple weather degradations under non-uniform illumination. To facilitate the research, we construct AllWeatherNight, a dataset for restoring nighttime images affected by multiple adverse weather effects and flares (see Fig. 2). Specifically, we collect diverse nighttime images from multiple sources and design an illumination-aware degradation generation method to faithfully emulate degradations in real-world nighttime photos, yielding images characterized by the realistic intertwined effects of uneven illumination and multiple adverse weather degradations. Consequently, training with our synthesized images enables models to generalize better in real-world nighttime scenarios.
@@
To effectively handle the intertwined degradations caused by artificial lighting and multiple adverse weather effects at night, we present ClearNight, the first unified framework tailored for multi-weather nighttime image restoration. We first leverage Retinex-based dual prior guidance to explicit disentangle the lighting and texture information within degraded images. Specifically, the decoupled illumination prior guides the model to focus on the uneven lit regions for effective restoration in nighttime scenes, while the reflectance prior enhances texture representations to mitigate weather-induced degradations and recover clear background details.
@@
In order to better represent complex degradations consisting of multiple adverse weather conditions, ClearNight employs a specificity-commonality dual-branched architecture, where the specificity branch is dynamically constructed and synergizes with the commonality branch. The dynamic specificity branch contains diverse sub-networks, formed by selecting different combinations of candidate units (losslb; specific; jia2024mos2; llm2024). To further enhance the capabilities of handling multi-weather degradations, we associate these candidate units with specific weather types using an auxiliary weather instructor, which identifies weather degradations and implicitly guides their dynamic allocation. Thus, the dynamic allocator becomes weather-aware and is encouraged to consistently select appropriate units for the same types of weather conditions.
Figure 2: Real-world and synthetic samples in our AllWeatherNight dataset. The synthetic images effectively simulate real-world nighttime scenes with various degradations.
@@
  • We propose ClearNight, the first unified framework for multi-weather nighttime image restoration, which integrates Retinex-based dual prior guidance and weather-aware dynamic specificity-commonality collaboration, tailored to address challenging uneven lighting and diverse weather effects entangled in nighttime scenes.
@@
  • Our ClearNight effectively removes various degradations and outperforms state-of-the-art approaches on both synthetic and real-world adverse weather nighttime images.
Related Work
@@
Adverse Weather Nighttime Image Restoration. Early works (NICIP2014; NCVPR2017) typically relied on physical priors and statistical assumptions, limiting their effectiveness and robustness in handling real-world scenarios. To address the limitations, several data-driven methods have been developed, achieving impressive results in restoring nighttime images degraded by haze (NICIP2012; NACM20232; NCVPR2024; NTIP2022; retinex7; CVPRW2022; NAAAI2025), rain (NICCV2023; NAAAI2024) and snow (SICCV2023). However, existing works focus on single-type degradations and overlook the fact that real-world nighttime scenes often involve multiple simultaneous degradations caused by adverse weather and uneven lighting.
@@
Multi-Weather Image Restoration. Multi-weather image restoration aims to restore various weather-degraded scenes using a single model (UMWECCV2022; FewECCV2024; LCVPR2024; MPPCVPR2024; LAAAI2024; RAAAI2024; DAAAI2025; UICLR2025; TSD2025). Recently, numerous Transformer-based approaches are developed, which investigate weather queries (TransCVPR2022), adverse weather pixel categorization (HistogramECCV2024), texture-guided appearance flow (CPCVPR2023), codebook prior fusion (AWICCV2023), and etc. Moreover, several recent works (WeatherDiff2023; TTTECCV2024; SCVPR2024) leverage remarkable generative capabilities of diffusion models for restoration performance boosting. To the best of our knowledge, all the previous methods focus on daytime scenes and ignore the entanglement of multiple weather conditions with non-uniform lighting in nighttime scenes.
@@
Adverse Weather Datasets. Most previous datasets, like Outdoor-Rain (HCVPR2019), RainDS (GOCVPR2021), BID-II (BECCV2022), CDD-11 (OneRestoreECCV2024) and WeatherStream (WSCVPR2023), only focus on adverse weather image restoration in daytime scenarios. A few recent studies explore this task in nighttime scenes and build datasets for haze (e.g. Night/YellowHaze (NPCM2018), NHC/NHM/NHR (NACM2020), UNREAL-NH (NACM20232)), rain (e.g. GTA5 (NECCV2020), GTAV-NightRain (GTAV; NICCV2023), Raindrop Clarity (NRaindropECCV2024)) and snow (e.g. RVSD (SICCV2023)). However, each existing nighttime dataset only considers a single type of weather condition, ignoring the challenge of mixed weather conditions commonly found in the real world.
AllWeatherNight Dataset
Real-world nighttime scenes are often degraded by a complex interplay of co-occurring weather and flares, challenging models for faithful restoration and visibility enhancement. To address this gap, we introduce AllWeatherNight, a new large-scale dataset featuring nighttime images with coupled degradations from mixed weather and uneven lighting.
Figure 3: Visualization of four synthesized variants of complex rain scene, where Weather Only and Flare Only denote synthesis with illumination-aware weather degradation and flare, respectively. Ours involves both degradations.
Data Collection and Filtering. We repurpose images from BDD100K (BDD100K) and ExDark (Exdark) as ground truths. These candidate images undergo a two-step process: selecting high-quality nighttime samples using average brightness, average gradient and grayscale variance; and manually selecting the 1,000 diverse images from each dataset. We also collect 1,000 real-world nighttime images with various weather conditions from the Internet and existing datasets (RESIDE; SPA; snow100k; GAN; DDN; NACM2023; JORDER).
Figure 4: Distribution of the our AllWeatherNight dataset.
Figure 5: Overview of our ClearNight framework. ClearNight primarily comprises Retinex-based dual prior guidance as well as weather-aware dynamic specificity and commonality branches. The Retinex-based dual priors explicitly guide the network to focus on illumination regions and intrinsic textures. The weather-aware dynamic specificity branch adaptively accommodates various weather effects and collaborates with the commonality branch to effectively handle complex multi-weather scenes.
Illumination-Aware Degradation Generation. We observe that uneven lighting conditions in real-world nighttime scenes often interact with adverse weather degradations, yet this phenomenon is largely overlooked by prior works (NACM2020; NICCV2023). To synthesize more realistic nighttime images with weather effects, we introduce an illumination-aware degradation generation approach that first modulates the lighting effects and then superimposes an illumination-aware combination of weather degradations.
Given a clean nighttime image $\mathit{X}$, we generate flare-degraded image $\mathit{X}^{\mathrm{flare}}$ based on light regions:
$\mathit{X}^{\mathrm{flare}}=\alpha\cdot\mathit{X}+\beta\cdot(\mathit{L}\ast\mathit{K}^{\mathrm{APSF}}),$ (1)
where $\alpha$ controls clean image preservation. $\beta$ regulates flare blending and is adaptively set by the light source pixel ratio. $\mathit{L}$ is a light source map devised via thresholding and alpha matting refinement (alpha). $\mathit{K}^{\mathrm{APSF}}$ is a 2D kernel from the atmospheric point spread function (NACM2023), and $\ast$ is the convolution operator.
Subsequently, building upon $\mathit{X}^{\mathrm{flare}}$, the illumination-aware weather-degraded image $\mathit{X}^{\mathrm{d}}$ is synthesized by:
$\mathit{X}^{\mathrm{d}}=\mathit{X}^{\mathrm{flare}}+\sum_{e\in\mathcal{E}}\omega_{e}\cdot\mathcal{F}^{\mathrm{G}}_{e}(\mathit{X}^{\mathrm{flare}}),$ (2)
where $\mathcal{E}$ is the set of selected weather effects for generating a specific degraded image, which is a subset of the universal set of all possible effects, i.e., $\mathcal{E}\subseteq\{\mathrm{H},\mathrm{RS},\mathrm{RD},\mathrm{S}\}$ (Haze, Rain Streak, Raindrop, and Snow). We devise a weight map $\omega_{e}$ to better simulate weather degradations under uneven lighting, where $\omega_{e\neq\mathrm{RD}}$ is set to the Retinex decomposed illumination map, while $\omega_{e=\mathrm{RD}}$ is empirically set to 1, as raindrops are primarily affected by local background rather than distant lighting. $\mathcal{F}^{\mathrm{G}}_{e}(\cdot)$ represents corresponding weather generation functions. We simulate adverse weather using established models (HCVPR2019; Raindrop; snow100k; csd). As shown in Fig. 3, unlike simplistic, spatially uniform synthesis, illumination-aware method generates more natural degradations.
Dataset Statistics. Overall, we synthesize 8,000 nighttime training images, covering multiple weather with varying scales, directions, patterns and intensities. As summarized in Fig. 4, the test dataset comprises synthetic and real-world subsets, each containing 1,000 images for model evaluation.
ClearNight
ClearNight is a unified nighttime image restoration framework designed to simultaneously remove multiple weather degradations. As depicted in Fig. 5, ClearNight integrates Retinex-based dual prior guidance and weather-aware dynamic specificity-commonality collaboration. The former explicitly decouples uneven lighting and textures to guide the network in recovering clear images with natural lighting and rich background details. The latter effectively captures the unique and shared characteristics of diverse weather conditions, enabling powerful multi-weather image restoration.
Retinex-Based Dual Prior Guidance
Decoupling illumination and texture information is critical for nighttime image restoration, as it allows the model to separately handle non-uniform lighting and adverse weather degradations. Retinex theory (retinex1) provides a classic physical model for this decomposition, formulating the input degraded image $\mathit{X}^{\mathrm{d}}$ as:
$\mathit{X}^{\mathrm{d}}=\mathit{R}^{\mathrm{d}}\cdot\mathit{I}^{\mathrm{d}},$ (3)
where $\mathit{R}^{\mathrm{d}}$ and $\mathit{I}^{\mathrm{d}}$ are the reflectance and illumination components, respectively. The decomposed components $\mathit{p}^{\mathrm{rtx}}\in\{\mathit{I}^{\mathrm{d}},\mathit{R}^{\mathrm{d}}\}$ are then fed into shared-weight multi-scale prior extraction units (MPE):
$\mathit{p}_{1},\mathit{p}_{2},\mathit{p}_{3}=\mathcal{F}^{\mathrm{MPE}}(\mathit{p}^{\mathrm{rtx}})\quad\textrm{with}\quad\mathit{p}_{n}\in\{\mathit{i}_{n},\mathit{r}_{n}\},$ (4)
where $\mathit{p}_{n}$ denotes the illumination or reflectance prior at the $n$-th scale, with $n\in\{1,2,3\}$. Within each MPE unit, a dilated convolution first projects the $\mathit{p}^{\mathit{rtx}}$ into three scales. Then, resulting multi-scale features are interactively fused to produce the final illumination/reflectance priors $\mathit{p}_{n}$.
More specifically, the illumination priors $\mathit{i}_{n}$ are successively injected into the Transformer blocks (TFBs) of the first three stages, which guide the network to focus on the uneven lighting regions in nighttime images, thereby facilitating the handling of lighting-influenced weather degradations. As the reflectance priors $\mathit{r}_{n}$ contain rich intrinsic textures that not only capture background details but also reveal degradation types, we incorporate them into each weather-aware dynamic selection block (WDS) to enhance weather type discrimination and improve multi-weather restoration.
Weather-Aware Dynamic Specificity-Commonality Collaboration
Dynamic Specificity and Commonality Synergy. As different weather degradations (e.g., snow and rain streaks) exhibit both shared and unique patterns, we introduce a synergistic design of dynamic specificity and commonality to effectively model complex multi-weather degradations. The commonality branch consists of sequential Transformer blocks (TFBs), while the specificity branch incorporates multiple weather-aware dynamic selection blocks (WDS) as residuals for each stage of TFBs.
As shown in the right of Fig. 5, each WDS consists of an encoder-decoder structure and a dynamic selection module, which jointly process the merged features from two branches and the reflectance prior. The encoder-decoder structure, inspired by (specific), learns compact representations, while the dynamic selection module constructs input-tailored sub-networks by sparsely selecting candidate units ($\mathcal{F}^{\mathrm{U}}_{k}(\cdot)$), where $k$ is the unit index.
Besides the candidate units, the dynamic selection module includes a gating ($\mathcal{F}^{\mathrm{W}}(\cdot)$) that computes the importance weights assigned to each candidate unit, and a router ($\mathcal{F}^{\mathrm{R}}(\cdot)$) that estimates the probability of each unit being selected. Given the input feature $\bar{\mathit{f}}_{a}$, the output of the $a$-th dynamic selection module can be calculated by
$\mathit{\bar{M}}_{a}=\sum_{k\in\mathcal{T}}\mathcal{F}^{\mathrm{W}}(\bar{\mathit{f}}_{a})_{k}\cdot\mathcal{F}^{\mathrm{U}}_{k}(\bar{\mathit{f}}_{a}),$ (5)
where $\mathcal{T}=\mathrm{TopK}(\mathcal{F}^{\mathrm{R}}(\bar{\mathit{f}}_{a}))$ denotes the set of selected top-$K$ candidate unit indices, and $\mathcal{F}^{\mathrm{W}}(\bar{\mathit{f}}_{a})_{k}$ is the importance weight for the $k$-th unit.
Figure 6: Comparison of 4 synthesized variants of nighttime images, where W and F indicate synthesis with illumination-aware weather degradation and flare effects respectively.
Dynamic Weather Degradation Modeling. Relying solely on visual content would limit the ability to capture the correlations and distinctive characteristics of different weather types, thereby hindering the module’s effectiveness in complex multi-weather scenarios. To better associate distinct weather types with designated candidate units, we reorganize the dynamic selection module and introduce a new component, the weather instructor ($\mathcal{F}^{\mathrm{WI}}(\cdot)$).
In the new design, features first pass through a knowledge-sharing unit ($\mathcal{F}^{\mathrm{KSU}}(\cdot)$), which is composed of shared linear layers preceding the dynamic allocator. The output of the knowledge-sharing unit is then utilized by the weather instructor, which classifies degradations and learns weather-specific prototypes to aggregate features from the same weather conditions:
$\mathit{y}_{a}=\mathcal{F}^{\mathrm{WI}}(\bar{\mathit{f}}_{a})\quad\textrm{with}\quad\bar{\mathit{f}}_{a}=\mathcal{F}^{\mathrm{KSU}}(\mathit{f}^{\prime}_{a}),$ (6)
where $\mathit{f}^{\prime}_{a}$ is the output of the encoder, and $\mathit{y}_{a}$ is the weather type predicted by the $a$-th dynamic selection module.
The weather instructor performs multi-label classification using binary cross-entropy loss $\mathcal{L}_{\mathrm{bce}}$, guiding the prototypes to attract features of the corresponding weather.
@@
Figure 8: Qualitative results on AllWeatherNight synthetic and real-world testing subset. Please zoom in for more details.